OpenStudy (anonymous):

Derivative of the Inverse of a Function: Let f(2)=6, f'(2)4, and h be the inverse of f. Find h'(6).

OpenStudy (anonymous):

\[(f^{-1})'(x)=\frac{1}{f'(f^{-1}(x))}\]

OpenStudy (anonymous):

and so \[(f^{-1})'(6)=\frac{1}{f'(f^{-1}(6))}\] you have all the numbers you need to compute

OpenStudy (31356):

@satellite73 help please?

OpenStudy (31356):

This is the last question.

OpenStudy (31356):

Is the answer -1d+d^2?

OpenStudy (anonymous):

clear or no? \(f(2)=6\) so \(f^{-1}(6)=2\) and \(f'(2)=4\) you get \(\frac{1}{4}\)

OpenStudy (anonymous):

hold on dear, let me go back and see

OpenStudy (31356):

Okay

OpenStudy (anonymous):

How do you know that \[f^-1 (6)= 2\]

OpenStudy (anonymous):

you know it because you are told that \(f(2)=6\) so necessarily \(f^{-1}(6)=2\) right?

OpenStudy (anonymous):

ohhh!!! okay!